OpenStudy (anonymous):

is 4x-2y=8 and y=2x-4 a solution or infinite

OpenStudy (anonymous):

i got infinite

OpenStudy (anonymous):

dear they are same equation : y=2x-4 \[\rightarrow 2x-y=4\rightarrow (2x-y=4)*2\rightarrow4x-2y=8 \] they are same equation so it's have infinitive answer